A waste disposal company is considering the replacement of one of its aging trucks. The key parameters of the three trucks under scrutiny are rovided below. Parameters Delta Epsilon Zeta 1. Initial Cost 250,000 375,000 450,000 230,000 at EOY1 2. Revenues increasing by 3.000 annually 2.5% annually thereafter 140,000 at EOY1 decreasing by 2,000 annually thereafter 195,000 at EOY1 235,000 at EOY1 increasing by | decreasing by 1% annually thereafter thereafter 125,000 at EOY1EOY1-EOY4 125,000; 3. Operating decreasing by EOY5-EOY8 135,000 2% annually costs | EOY9-EOY12-170,000 EOY13-EOY16 190,000 thereafter 4. End-of-life salvage value ($ -20,000 7,000 -20,000 5. Useful life 4 ears EOY End-of-year Industry Standard - 4 years MARR-10% . 13. The best truck based on the simple payback method is 14. Deltas benefit/cost (B/C) ratio (second decimal; no rounding) is 15. Zetas benefit/cost (B/C) ratio (second decimal; no rounding) is 16. The incremental B/C ratio (second decimal; no rounding) between the a) Delta; b) Epsilon; c) Zeta a) 0.98; b) 1.03; c) 1.08; d) 1.10 a) 0.99; b) 1.02; c) 1.06; d) 1.09 Delta and Epsilon trucks IS a) 0.90; b) 0.99; c) 1.03; d) 1.08

DISCRETE CASH FLOWS AND DISCRETE COMPOUNDING 10.00 % DISCRETE RATE OF INTEREST 0.1000 0.1200 if k İzk 0.9091 1.8182 .1000 0.9091 1000 0.90911.00001.00000.0000 0.0000 1.21000.8264 0.5762 1.7355 0.4762 2.1000 0.47620.8264 1.33100.7513 0.40212.4869 0.30213.31000.9366 2.3291 1.4641 0.6830 0.3155 3.16990.2155 4.6410 1.38124.37816.4100 0.0000 0.9091 1.00001.8347 3.10002.7772 4 4.5455 5.4545 6.3636 7.2727 8.1818 9.0909 1.7716 0.5645 0.2296 4.3553 0.1296 7.7156 2.2236 9.6842 17.1561 1.94870.5132 0.20544.8684 0.10549.4872 2.6216 12.763124.87176.7215 2.14360.4665 0.18745.33490.0874 11.43593.0045 16.0287 34.35897.7528 2.35790.4241 0.1736 5.75900.0736 13.57953.372419.4215 45.79488.8028 2.59370.38550.16276.1446 0.0627 15.93743.7255 22.8913 59.3742 9.8719 2.85310.35050.15406.49510.0540 18.5312 4.0641 26.3963 75.3117 10.9605 10.0000 3.1384 0.3186 0.1468 6.8137 0.0468 21.3843 4.3884 29.9012 93.8428 12.0689 10.9091 3.45230.28970.14087.10340.040824.5227 4.6988 33.3772 115.227113.197411.8182 3.7975 0.2633 0.1357 7.3667 0.0357 27.9750 4.9955 36.8005 139.7498 14.3465 12.7273 4.17720.2394 0.1315 7.6061 0.0315 31.7725 5.2789 40.1520 167.7248 5.516413.6364 100 12 133 14 15 100 8 7.8237 0.0 18 10 200 5.05450.1978 0.12478.0216 0.024740.54475.807146.5819 235.447017.9205 15.4545 5.55990.1799 0.1219 8.2014 0.0219 45.5992 6.0526 49.6395 275.9917 19.1554 16.3636 6.1159| 0.1635| 0.1195| 8.3649| 0.0195| 51.1591| б.2861| 52.5827 321.5909| 20.4128| 17.2727 6.7275 0.1486 0.11758.51360.017557.2750 6.508155.4069372.7500 21.693018.1818

Economics practice questions (please show your work and the answer could always be a 'none of the above' answer in addition to the options given)
Answer #1

13.

Payback period is the time (usually in years) required by a project to recover the invested money.

Formula = Total Investment ÷ (Net) Annual Cash Inflows

Calculating Cash Infow for the following:

Cash Inflows less Cash Outflows = Revenues less Operating Costs

Note: Salvage value is after the operating period, hence ignored.

Delta : 230,000 increasing at 2.5% for 4 years is 253,877

less

140,000 reduced by 2000 each year for 4 years is 134,000

equals 119,877

Eplison: 195,000 increasing at 3,000 each year for 8 years is 207,000

less

125,000 decreasing at 2% for 8 years is 108,515

equals 11,362

Zeta: 235,000 decreasing at 1% for 16 years is 200,093

less

Average of 4 years 155,000

equals 45,093

Payback Period for

Delta = 250,000 ÷ 119,877 = 2.08 years

Eplison = 375,000 ÷ 11,362 = 33 years

Zeta = 450,000 ÷ 45,093 = 9.97 years

Best truck is Delta.

14.

BCR or Benefit to Cost Ratio gives an idea about the feasibility of a project.

BCR of Delta = NPV of Benefits ÷ Net of Investment including salvage value

= 253,877 ÷ (250,000 + (-20,000))

= 1.10 Option (d)

15.

BCR of Zeta = 200,093 ÷ 430,000 = 0.46 None of the above

16.

Incremental BC Ratio = .75 (Sum of Benefits divided by Initial Costs plus Salvage Vales)

Hence, none of the above.